# arXiv:2603.03264v1[astro-ph.SR] 03 Mar 2026

### 03 Mar 2026 v1

@@ 5.3 Model-driven $\delta T$ variations
We applied the model-driven approach used to derive temperature variation (see Sec. 4.3) to the observed spectra considered in our study. For EV Lac (Gl 873) and DS Leo (Gl 410), we find significant correlations between our derived $\delta T$ and the data-driven estimates obtained with the LBL (Artigau et al., 2024) and published by Cristofari et al. (2025a), with Pearson correlation coefficients of 0.95 and 0.93, respectively (see Fig. 9). A moderate correlation is also observed for Barnard’s star, with a Pearson coefficient of 0.59 (see Fig. 9). We note the smaller range of the variations relative to error bars for this star. These results demonstrate the ability models have to reproduce relative temperature variations in spite of the significant discrepancies between models and observations (see, e.g., Cristofari et al. 2022b, a). We note that in a similar experiment, Artigau et al. (2024) found that model-driven variations led to a damping of the retrieved $\delta T$. Details in implementation and normalization, along with differences between the models used could explain our ability to retrieve amplitudes comparable to those obtained with the data-driven approach. Our results further support previous findings (Bellotti et al., 2025) where synthetic spectra were used to estimate temperature and $\langle B\rangle$ variations from spectra recorded with SPIRou.
@@
From the spectra recorded with SPIRou, we observe strong anti-correlations between our $\delta T$ and $\delta\langle B\rangle$ estimates. We apply orthogonal distance regressions222implementation relying on Scipy (Virtanen et al., 2020) (ODRs) to estimate the properties of the linear relationships between $\delta T$ and $\delta\langle B\rangle$ (see Fig. 6 and 10). For EV Lac, DS Leo and Barnard’s star, the fits yield slopes of $18.3\pm 0.6$, $54.2\pm 1.8$$\rm K\,kG^{-1}$, and $24.7\pm 1.0$$\rm K\,kG^{-1}$, consistent with those reported by Cristofari et al. (2025a, $19.8\pm 0.5$, $60.4\pm 2.5$ and $18.8\pm 0.8$$\rm K\,kG^{-1}$, respectively). We note that the ODR is very sensitive to the relative scaling of the error bars of both axes, which can have an impact of several $\rm K\,kG^{-1}$ in our case.
@@
From the data recorded with Narval/ESPaDOnS, we observe similar anticorrelations between $\delta T$ and $\delta\langle B\rangle$, with significantly larger scatters, yielding slopes of $22.3\pm 3.3$ and $61.7\pm 6.8$$\rm K\,kG^{-1}$ for EV Lac and DS Leo, respectively. The larger scatters in the results could be attributed to the impact of magnetic fields on $\delta T$ estimates (see Sec 4.3), the sparsity of the observations obtained with Narval/ESPaDOnS, or the higher contrast in the optical where cool spots could lead to larger temperature variations which may not be strictly anticorrelated with $\delta B$.
@@
We note that while our simulations suggest that the impact of magnetic fields on the derived $\delta T$ is more significant for the Narval/ESPaDOnS domain, it could still have a non-negligible impact on the estimates obtained from spectra recorded in the nIR, which could lead to steeper slopes and larger scattering. A natural next step will be to use the $\delta\langle B\rangle$ estimates to obtain magnetic field dependent $\partial\@vec{A}/\partial T$ (see equation 4), and rely on these to refine the $\delta T$ measurements. We leave the implementation of this process for a future study.
@@ 5.3.1 Comparison to activity indicators
From Narval and ESPaDOnS spectra, we measured the S-index, commonly used as an activity proxy, to compare it to our $\delta\langle B\rangle$ estimates. We follow the definition of the S-index of Vaughan et al. (1978):
$\text{S-index}=\frac{aF_{H}+bF_{K}}{cF_{R}+dF_{V}}+e$
where $F_{H}$ and $F_{K}$ are the fluxes taken in $1.09\,\AA$ triangular bands centered on $3968.470$ and $3933.661\,\AA$ , respectively . $F_{R}$ and $F_{V}$ are the fluxes computed in $20\,\AA$ rectangular bands centered on $3901$ and $4001\,\AA$ , respectively (see Fig. 17). The coefficients $a$, $b$, $c$, $d$ and $e$ allow one to calibrate the index to the Mount Wilson scale and were taken from Marsden et al. (2014). Flux uncertainties were propagated to estimate error bars on each band’s flux, which were in turn used to define error bars on the S-index measurements.
We remind that $\delta\langle B\rangle$ is directly linked to $\langle B\rangle$ such that $\langle B\rangle=\delta\langle B\rangle+\langle B\rangle_{\rm ref}$, with $\langle B\rangle_{\rm ref}$ the value associated with the reference spectrum (see equation 4). For easier comparison with the literature, in Fig. 7, we compute $\langle B\rangle$ from $\delta\langle B\rangle$ taking the typical $\langle B\rangle_{\rm ref}$ values from Cristofari et al. (2025a) for EV Lac and DS Leo ($4.54$ and $0.79$ kG, respectively). No clear correlation is observed between the measured S-index and the magnetic field strength variation derived with our approach for EV Lac or DS Leo (Fig. 7). We note that the long baseline of the observations ($\sim 10$ yrs) could encompass activity regime changes that could break the activity-$\langle B\rangle$ relationship that one would expect for an activity driven by quiescent heating. For both stars, we divided the observations in segments of closely spaced observations and compared the activity indicators to $\delta\langle B\rangle$ in these. A tentative correlation (with a Pearson coefficient of 0.57) can be observed for DS Leo if one only considers data points between BJD 2454400 and 2455700. These dates encompass observation campaigns where the magnetic field strength is at its lowest. These results further suggest that changes in activity regimes could throw off the relation between the S-index and average surface magnetic field strength on long observation baselines. Recently, Hahlin et al. (2026) reported correlations between the S-index and $\langle B\rangle$ for a few hotter stars ($T_{\rm eff}$ ranging from $\sim$ 4350 to 5400 K), suggesting that such relation may be harder to recover for M dwarfs than for Sun-like stars.
@@
Figure 7: S-index compared to $\langle B\rangle$ for EV Lac (top) and DS Leo (bottom). $\langle B\rangle$ estimates were obtained from $\delta\langle B\rangle$ by adding typical median values from Cristofari et al. (2025a), of 4.54 and 0.79 kG for EV Lac and DS Leo, respectively. For DS Leo, points marked with a cyan circle correspond to observations recorded between BJD 2454400 and 2455700 (see Fig. 5).
@@ 6 Discussion and conclusions
With this paper, we study the temporal behavior of magnetic fields from long-term spectroscopic monitoring of cool stars (see, e.g., Donati et al., 2023a; Cristofari et al., 2025a). We explore a simple method to estimate relative variations of small-scale magnetic fields ($\delta\langle B\rangle$) from high-resolution spectra. The proposed method does not rely on fits of synthetic spectra to individual observations, providing extremely fast estimates in comparison.
@@
Studies relying on fits of synthetic spectra to observations (e.g., Reiners et al., 2022; Cristofari et al., 2025a) are inherently limited by the quality of the fits, with systematics between modeled and observed line shapes leading to uncertainties and potential biases in the results. The method proposed in this paper attempts to mitigate these effects by only studying changes in line shapes relative to a reference spectrum, obtained from the observations themselves. While the fundamental idea of studying spectral feature variations compared to a reference spectrum is not new, with notable applications for radial velocity (Artigau et al., 2022) and surface temperature variations (Artigau et al., 2024), this is to our knowledge the first such attempt to derive relative small-scale magnetic field variations ($\delta\langle B\rangle$).
@@
With this work, we show that our model-informed process applied to observations can provide reliable information on $\delta\langle B\rangle$ from series of high-resolution spectra. Our approach relied on synthetic spectra computed with ZeeTurbo (Cristofari et al., 2023a), and was applied to high-resolution near-infrared and optical spectra recorded with SPIRou, Narval and ESPaDOnS. Our results were compared to magnetic field measurements previously obtained for a few strongly-magnetic targets (Cristofari et al., 2025a), yielding excellent agreement. This underscores the reliability of the process, which by construction allows one to obtain $\delta\langle B\rangle$ from a linear fit, instead of expensive Markov-Chain Monte-Carlo sampling (Cristofari et al., 2025a). Consequently, the proposed method allows one to treat hundreds of spectra in seconds, making it a diagnostic tool viable for systematic application to spectra recorded in the context of the large surveys to come.
@@
The analysis of Narval and ESPaDOnS spectra provide new constraints on the long-term variations of the small-scale magnetic fields for EV Lac and DS Leo. Quasi-Periodic Gaussian Process fits to the series of $\delta\langle B\rangle$ measurements show a clear modulation consistent with the rotation period of the stars. We note that this is achieved in spite of the relative sparsity of the data over a 10 years period, compared the dense survey carried out with SPIRou over 3 years.
@@
Through simulations and analysis of the observed data, we investigated the impact that assumptions on atmospheric properties have on the results, and found that our method is relatively insensitive to small changes in the adopted parameters. We also found that significant offset in the assumed atmospheric properties leads to a damping of the derived $\delta\langle B\rangle$. Consequently, even with an incorrect stellar characterization, our method provides valuable information on the temporal behavior of the magnetic fields, with a damped amplitude. Our simulations also underscore the importance of careful relative normalization, as short-window median filters tend to impact the shape of broad absorption features, with significant impact on both $\delta\langle B\rangle$ and temperature variation estimates ($\delta T$). The proposed method allows one to retrieve extremely fast estimates of the absolute surface magnetic field strength ($\langle B\rangle$) for a series of spectra, provided that the value associated with the reference spectrum is known. Such an initial estimate can be obtained with the methods commonly used to derive field strengths (e.g., Reiners et al., 2022; Cristofari et al., 2023b; Kochukhov et al., 2024).
@@
We implemented a model-driven process following Artigau et al. (2024) to derive relative temperature variations ($\delta T$) from the observed spectra. For EV Lac and DS Leo, we find excellent agreement between our results and the data-driven $\delta T$ obtained in the framework of the LBL code by (Artigau et al., 2024). Similarly to Cristofari et al. (2025a), we find significant anti-correlation between $\delta T$ and $\delta\langle B\rangle$, suggesting that the observed magnetic fields are directly linked to the presence of cooler spots at the stellar surface. Relying on spectra recorded with ESPaDOnS and Narval, we see clear anti-correlations between our retrieved $\delta T$ and $\delta\langle B\rangle$ for EV Lac and DS Leo. We obtain similar slopes between $\delta T$ and $\delta\langle B\rangle$ for the optical and nIR domains, although with a significantly higher scatter in the optical domain. Several competing factors could explain these results, including the possibly higher sensitivity to temperature variations in the optical, caused by higher contrast between spots and the photosphere (Berdyugina, 2005). One identified cause, however, could be the influence of magnetic fields on the $\delta T$ estimates. Simulations revealed that magnetic field variations can lead to an increased scatter and larger amplitudes in the derived $\delta T$ (Fig. 2), in particular for regions containing a large number of magnetically-sensitive transitions, which is consistent with the larger scatter observed for ESPaDOnS and Narval data. We note that a natural next step could be to rely on the $\delta\langle B\rangle$ measurements to refine $\delta T$ estimates. Using $\delta\langle B\rangle$, one could compute magnetic-field dependent $\partial\@vec{A}/\partial T$ (equation 4), and use these to obtain more accurate $\delta T$. We leave the detailed implementation of this process to a subsequent study. We note that temperature variations can also have a small impact on the derived $\delta\langle B\rangle$, although these appear small in light of the current measurements precision. The next generation of instruments will provide spectra with much higher quality, and the implementation of temperature dependent magnetic models, mimicking the impact of cool magnetic spots at the stellar surface, will enable a more detailed modeling of the stellar spectra.
@@
The current method will benefit from improvement in the models used to obtain the relative variation of pixels with magnetic field strength. The models used in this work were computed under the assumption that only atoms are sensitive to magnetic fields, due to the lack of known Landé factors for most magnetically-sensitive molecules. While the impact of magnetically-sensitive molecules on the results can be mitigated by pixel rejection strategies and avoiding dense molecular bands, including the impact of magnetic fields on sensitive molecules (such as TiO) will allow to increase the precision of the proposed method.
The proposed method can have significant benefits for the correction of radial velocity curves in the context of exoplanet searches. Recent studies suggested that $\langle B\rangle$ can be an excellent indicator of activity jitters in radial velocity curves (Haywood et al., 2016, 2022). The ability to obtain fast and reliable constraints from the same data used to measure radial velocities opens the door to new strategies for activity jitter correction in order to extract planetary signals from the recorded data.

### 26 Mar 2026 v2

@@ 5.3 Model-driven $\delta T$ variations
We applied the model-driven approach used to derive temperature variation (see Sect. 4.3) to the observed spectra considered in our study. For EV Lac (Gl 873) and DS Leo (Gl 410), we find significant correlations between our derived $\delta T$ and the data-driven estimates obtained with the LBL (Artigau et al., 2024) and published by Cristofari et al. (2025), with Pearson correlation coefficients of 0.95 and 0.93, respectively (see Fig. 9). A moderate correlation is also observed for Barnard’s star, with a Pearson coefficient of 0.59 (see Fig. 9). We note the smaller range of the variations relative to error bars for this star. These results demonstrate the ability models have to reproduce relative temperature variations in spite of the significant discrepancies between models and observations (see, e.g., Cristofari et al. 2022b, a). We note that in a similar experiment, Artigau et al. (2024) found that model-driven variations led to a damping of the retrieved $\delta T$. Details in implementation and normalization, along with differences between the models used could explain our ability to retrieve amplitudes comparable to those obtained with the data-driven approach. Our results further support previous findings (Bellotti et al., 2025) where synthetic spectra were used to estimate temperature and $\langle B\rangle$ variations from spectra recorded with SPIRou.
@@
From the spectra recorded with SPIRou, we observed strong anti-correlations between our $\delta T$ and $\delta\langle B\rangle$ estimates. We applied orthogonal distance regressions [^2] Implementation relying on Scipy (Virtanen et al., 2020) (ODRs) to estimate the properties of the linear relationships between $\delta T$ and $\delta\langle B\rangle$ (see Fig. 6 and 10). For EV Lac, DS Leo, and Barnard’s star, the fits yield slopes of $18.3\pm 0.6$, $54.2\pm 1.8$$\rm K\,kG^{-1}$, and $24.7\pm 1.0$$\rm K\,kG^{-1}$, consistent with those reported by Cristofari et al. (2025, $19.8\pm 0.5$, $60.4\pm 2.5,$ and $18.8\pm 0.8$$\rm K\,kG^{-1}$, respectively). We note that the ODR is very sensitive to the relative scaling of the error bars of both axes, which can have an impact of several $\rm K\,kG^{-1}$ in our case.
@@
From the data recorded with Narval/ESPaDOnS, we can observe similar anticorrelations between $\delta T$ and $\delta\langle B\rangle$, with significantly larger scatters, yielding slopes of $22.3\pm 3.3$ and $61.7\pm 6.8$$\rm K\,kG^{-1}$ for EV Lac and DS Leo, respectively. The larger scatters in the results could be attributed to the impact of magnetic fields on $\delta T$ estimates (see Sec 4.3), the sparsity of the observations obtained with Narval/ESPaDOnS, or the higher contrast in the optical where cool spots could lead to larger temperature variations which may not be strictly anticorrelated with $\delta B$.
@@
We note that while our simulations suggest that the impact of magnetic fields on the derived $\delta T$ is more significant for the Narval/ESPaDOnS domain, it could still have a non-negligible impact on the estimates obtained from spectra recorded in the nIR, which could lead to steeper slopes and larger scattering. A natural next step would be to use the $\delta\langle B\rangle$ estimates to obtain magnetic field-dependent $\partial\@vec{A}/\partial T$ values (see Eq. 4) and rely on these to refine the $\delta T$ measurements. We leave the implementation of this process to a future study.
@@ 5.3.1 Comparison to activity indicators
From Narval and ESPaDOnS spectra, we measured the S-index, commonly used as an activity proxy, to compare it to our $\delta\langle B\rangle$ estimates. We followed the definition of the S-index from Vaughan et al. (1978):
$\text{S-index}=\frac{aF_{H}+bF_{K}}{cF_{R}+dF_{V}}+e$ (6)
where $F_{H}$ and $F_{K}$ are the fluxes taken in $1.09\,\AA$ triangular bands centered on $3968.470$ and $3933.661\,\AA$ , respectively. $F_{R}$ and $F_{V}$ are the fluxes computed in $20\,\AA$ rectangular bands centered on $3901$ and $4001\,\AA$ , respectively (see Fig. 17). The coefficients $a$, $b$, $c$, $d$, and $e$, which allow us to calibrate the index to the Mount Wilson scale, were taken from Marsden et al. (2014). The flux uncertainties were propagated to estimate the error bars on each band’s flux, which were, in turn, used to define error bars on the S-index measurements.
We recall here that $\delta\langle B\rangle$ is directly linked to $\langle B\rangle$ such that $\langle B\rangle=\delta\langle B\rangle+\langle B\rangle_{\rm ref}$, with $\langle B\rangle_{\rm ref}$ the value associated with the reference spectrum (see Eq. 4). For an easier comparison with the literature, in Fig. 7, we computed $\langle B\rangle$ from $\delta\langle B\rangle,$ taking the typical $\langle B\rangle_{\rm ref}$ values from Cristofari et al. (2025) for EV Lac and DS Leo ($4.54$ and $0.79$ kG, respectively). No clear correlation was observed between the measured S-index and the magnetic field strength variation derived with our approach for EV Lac or DS Leo (Fig. 7). We note that the long baseline of the observations ($\sim 10$ yrs) could encompass activity regime changes that could break the activity-$\langle B\rangle$ relationship that would be expected for an activity driven by quiescent heating. For both stars, we divided the observations in segments of closely spaced observations and compared the activity indicators to $\delta\langle B\rangle$ in these. A tentative correlation (with a Pearson coefficient of 0.57) can be observed for DS Leo if we are only considering data points between BJD 2454400 and 2455700. These dates encompass observation campaigns where the magnetic field strength is at its lowest. These results further suggest that changes in activity regimes could throw off the relation between the S-index and average surface magnetic field strength on long observation baselines. Recently, Hahlin et al. (2026) reported correlations between the S-index and $\langle B\rangle$ for a few hotter stars ($T_{\rm eff}$ ranging from $\sim$ 4350 to 5400 K), suggesting that such a relation might be harder to recover for M dwarfs than it would be for Sun-like stars.
@@
Figure 7: S-index compared to $\langle B\rangle$ for EV Lac (top) and DS Leo (bottom). $\langle B\rangle$ estimates were obtained from $\delta\langle B\rangle$ by adding typical median values from Cristofari et al. (2025), of 4.54 and 0.79 kG for EV Lac and DS Leo, respectively. For DS Leo, points marked with a cyan circle correspond to observations recorded between BJD 2454400 and 2455700 (see Fig. 5).
@@ 6 Discussion and conclusions
In this paper, we study the temporal behavior of magnetic fields from long-term spectroscopic monitoring of cool stars (see, e.g., Donati et al., 2023a; Cristofari et al., 2025). We explore a simple method to estimate relative variations of small-scale magnetic fields ($\delta\langle B\rangle$) from high-resolution spectra. The proposed method does not rely on fits of synthetic spectra to individual observations, providing extremely fast estimates in comparison.
@@
Studies that rely on fits of synthetic spectra to observations (e.g., Reiners et al., 2022; Cristofari et al., 2025) are inherently limited by the quality of the fits, with systematics between modeled and observed line shapes leading to uncertainties and potential biases in the results. The method proposed in this paper attempts to mitigate these effects by only studying changes in line shapes relative to a reference spectrum, obtained from the observations themselves. While the fundamental idea of studying spectral feature variations compared to a reference spectrum is not new, with notable applications for radial velocity (Artigau et al., 2022) and surface temperature variations (Artigau et al., 2024), this is to our knowledge the first such attempt to derive relative small-scale magnetic field variations ($\delta\langle B\rangle$).
@@
With this work, we show that our model-informed process applied to observations can provide reliable information on $\delta\langle B\rangle$ from series of high-resolution spectra. Our approach relies on synthetic spectra computed with ZeeTurbo (Cristofari et al., 2023a), applied to high-resolution near-infrared and optical spectra recorded with SPIRou, Narval, and ESPaDOnS. Our results were compared to magnetic field measurements previously obtained for a few strongly magnetic targets (Cristofari et al., 2025), yielding an excellent agreement. This underscores the reliability of the process, which (by construction) allows us to obtain $\delta\langle B\rangle$ from a linear fit, instead of expensive MCMC sampling (Cristofari et al., 2025). Consequently, the proposed method enables the treatment of hundreds of spectra in seconds, making it a diagnostic tool that is viable for systematic application to spectra recorded in the context of large surveys to come.
@@
The analysis of Narval and ESPaDOnS spectra provide new constraints on the long-term variations of the small-scale magnetic fields for EV Lac and DS Leo. QPGP fits to the series of $\delta\langle B\rangle$ measurements show a clear modulation consistent with the rotation period of the stars. We note that this has been achieved in spite of the relative sparsity of the data over a ten-year period, compared the dense survey carried out with SPIRou over three years.
@@
Through simulations and analysis of the observed data, we investigated the impact that assumptions on atmospheric properties have on the results and found that our method is relatively resilient with respect to small changes in the adopted parameters. We also found that significant offset in the assumed atmospheric properties leads to a damping of the derived $\delta\langle B\rangle$. Consequently, even with an incorrect stellar characterization, our method provides valuable information on the temporal behavior of the magnetic fields, with a damped amplitude. Our simulations also underscore the importance of careful relative normalization, as short-window median filters tend to impact the shape of broad absorption features, with significant impacts on both $\delta\langle B\rangle$ and temperature variation estimates ($\delta T$). The proposed method allows for the retrieval of extremely fast estimates of the absolute surface magnetic field strength ($\langle B\rangle$) for a series of spectra, provided that the value associated with the reference spectrum is known. Such an initial estimate can be obtained with the methods commonly used to derive field strengths (e.g., Reiners et al., 2022; Cristofari et al., 2023b; Kochukhov et al., 2024).
@@
We implemented a model-driven process following Artigau et al. (2024) to derive relative temperature variations ($\delta T$) from the observed spectra. For EV Lac and DS Leo, we found an excellent agreement between our results and the data-driven $\delta T$ obtained in the framework of the LBL code by (Artigau et al., 2024). Similarly to Cristofari et al. (2025), we find significant anticorrelation between $\delta T$ and $\delta\langle B\rangle$, suggesting that the observed magnetic fields are directly linked to the presence of cooler spots at the stellar surface. Relying on spectra recorded with ESPaDOnS and Narval, we see clear anticorrelations between our retrieved $\delta T$ and $\delta\langle B\rangle$ for EV Lac and DS Leo. We obtained similar slopes between $\delta T$ and $\delta\langle B\rangle$ for the optical and nIR domains, although with a significantly higher scatter in the optical domain. Several competing factors could explain these results, including the possibly higher sensitivity to temperature variations in the optical, caused by higher contrast between spots and the photosphere (Berdyugina, 2005). One identified cause, however, could be the influence of magnetic fields on the $\delta T$ estimates. Simulations revealed that magnetic field variations can lead to an increased scatter and larger amplitudes in the derived $\delta T$ (Fig. 2), particularly for regions containing a large number of magnetically sensitive transitions, which is consistent with the larger scatter observed for ESPaDOnS and Narval data. We note that a natural next step could be to rely on the $\delta\langle B\rangle$ measurements to refine $\delta T$ estimates. Using $\delta\langle B\rangle$, we could compute magnetic-field dependent $\partial\@vec{A}/\partial T$ (Eq. 4) and used these to obtain more accurate $\delta T$. We leave the detailed implementation of this process to a subsequent study. We note that temperature variations can also have a small impact on the derived $\delta\langle B\rangle$, although these appear small in light of the current measurements precision. The next generation of instruments will provide spectra with much higher quality and the implementation of temperature dependent magnetic models, mimicking the impact of cool magnetic spots at the stellar surface, will enable a more detailed modeling of the stellar spectra.
@@
The current method will benefit from improvement in the models used to obtain the relative variation of pixels with magnetic field strength. The models used in this work were computed under the assumption that only atoms are sensitive to magnetic fields, due to the lack of known Landé factors for most magnetically sensitive molecules. While the impact of magnetically sensitive molecules on the results can be mitigated by pixel rejection strategies and avoiding dense molecular bands, including the impact of magnetic fields on sensitive molecules (such as TiO) will allow for the precision of the proposed method to be improved.
Finally, we find that the proposed method can have significant benefits for the correction of radial velocity curves in the context of exoplanet searches. Recent studies have suggested that $\langle B\rangle$ can be an excellent indicator of activity jitters in radial velocity curves (Haywood et al., 2016, 2022). The ability to obtain fast and reliable constraints from the same data used to measure RVs opens the door to new strategies for activity jitter correction for extracting planetary signals from the recorded data.